# RRWNet: Recursive Refinement Network for Effective Retinal Artery/Vein Segmentation and Classification

### v2

@@ 5.1 Datasets
Experiments were performed on the three publicly available datasets containing color fundus images with corresponding A/V annotations: RITE Hu et al., (2013), LES-AV Orlando et al., (2018), and HRF Budai et al., (2013). Fig. 5 shows some examples of color fundus images and their corresponding GT segmentation maps from the three datasets, while Table 1 provides an overview of the distribution of samples (pixels) across the different classes (background, artery, vein, crossing, uncertain) within each dataset. Further details on the datasets are provided below.
@@
The RITE [^1] https://medicine.uiowa.edu/eye/rite-dataset (accessed on 2023-12-15) dataset Hu et al., (2013) (also known as AV-DRIVE or DRIVE-AV in the literature) is an extension of the Digital Retinal Images for Vessel Extraction (DRIVE) dataset Staal et al., (2004). While DRIVE focuses on vessel segmentation, RITE incorporates additional manual GT segmentation maps for classifying arteries and veins. The dataset comprises the same 40 retinography of DRIVE (20 for training and 20 for testing). These images originate from 33 healthy patients and 7 patients with mild signs of DR. They are all centered on the macula and have a resolution of $768\times 584$ pixels, with a circular region of interest (ROI). RITE includes the original blood vessel segmentation maps from DRIVE, along with the pixel-level classification of these vessels into arteries, veins, crossings, and uncertain classes Hu et al., (2013). Crossings indicate areas where a vein and an artery overlap. The “uncertain” class is used for those vessels whose classification the experts have not been able to determine. Alternative A/V classification annotations for the DRIVE dataset were proposed by Qureshi et al., (2013). They manually segmented and classified the blood vessels into arteries, veins, and uncertain from the raw retinography images, independent of the existing DRIVE vessel segmentations. For artery-vein crossings, the classification was assigned based on the vessel closest to the surface of the retina. These alternative labels, henceforth referred to as Qureshi et al., were used in this work as additional “second expert” annotations.
@@
This dataset [^2] https://figshare.com/articles/dataset/LES-AV_dataset/11857698 (accessed on 2023-12-16) Orlando et al., (2018) consists of 22 retinography images, originating from both healthy patients (11) and patients with signs of glaucoma (11). Unlike RITE, LES-AV does not have a predefined split for training and testing. The images are centered on the optic disc and have a resolution of $1620\times 1444$ pixels (except one image with a resolution of $2196\times 1958$ pixels) and a circular ROI. Similarly to RITE, LES-AV includes GT segmentation maps for blood vessels, classified into arteries, veins, crossings, and uncertain regions. In this work, we employ LES-AV as an external dataset for cross-dataset evaluations to assess the generalization capabilities of the proposed method.
@@
The HRF dataset [^3] https://www5.cs.fau.de/research/data/fundus-images/ (accessed on 2023-12-16) Budai et al., (2013) is a collection of 45 high-resolution retinography images ($3504\times 2336$ pixels). The images are categorized into three groups: 15 images from healthy individuals, 15 from patients with diabetic retinopathy (DR), and 15 from glaucomatous patients. A/V classification annotations were provided by two different sources. Hemelings et al., (2019) [^4] https://github.com/rubenhx/av-segmentation (accessed on 2023-12-16) provided the original annotations used for the dataset. Then, Chen et al., (2022) [^5] https://github.com/o0t1ng0o/TW-GAN (accessed on 2023-12-16) introduced novel manual annotations to address inconsistencies in the annotations of Hemelings et al., (2019). The annotation procedures of these two works differ slightly in handling artery-vein crossings. While Chen et al., (2022) label crossings as a separate class (consistent with RITE and LES-AV), Hemelings et al., (2019) assign them to the uppermost vessel if known, or to the uncertain class otherwise. In this work, we primarily use the Chen et al., (2022) annotations for training and testing, while we use the Hemelings et al., (2019) annotations as “second expert” annotations. Following previous works Hemelings et al., (2019); Karlsson and Hardarson, (2022), we use the first five images in each category for testing and the remainder for training.
@@ 5.2 Experiments
We performed an ablation study to evaluate the individual and combined impact of the proposed RR module and the recursive refinement (multiple RR applications). For this end, we compared the following models: (1) U-Net: A standard U-Net architecture Ronneberger et al., (2015). This is equivalent to use only the Base subnetwork of RRWNet. (2) W-Net: Our proposed approach without recursive refinement (i.e., $K=1$). This isolates the effect of the recursive refinement. (3) RRU-Net: An architecture similar to Mosinska et al., (2018), that employs recursive refinement with a single encoder-decoder. This is equivalent to use only the RR subnetwork of RRWNet in a recursive manner while providing the input image. (4) RRWNetAll: Our full model (Base and RR subnetworks) refining all segmentation maps (arteries, veins, and BV). (5) RRWNet: Our full model refining only the A/V segmentation maps. All experiments were conducted on the RITE dataset. Statistical significance was assessed using the one-tailed Wilcoxon signed-rank test.
@@
We compared RRWNet with several state-of-the-art methods for A/V segmentation and classification on RITE, LES-AV and HRF datasets: Girard et al., (2019); Galdran et al., (2019); Ma et al., (2019); Hemelings et al., (2019); Kang et al., (2020); Morano et al., (2021); Chen et al., (2022); Galdran et al., (2022); Hatamizadeh et al., (2022); Karlsson and Hardarson, (2022); Hu et al., (2024). To comprehensively evaluate our approach, we used different evaluation protocols. In the first place, we employed the standard evaluation protocol in the field, focusing on the A/V classification accuracy for the intersection of predicted and GT vessels. Additionally, we performed a more in depth comparison with recent state-of-the-art methods that provided source code or continuous predicted segmentation maps for any of the datasets. In particular, following Morano et al., (2021) and Chen et al., (2022), we compared the methods in terms of A/V classification performance for all vessel pixels in the GT, as well as as A/V segmentation performance using and several threshold-independent and topological metrics. The methods included in the comparison were Morano et al., (2021), Chen et al., (2022) [^6] https://github.com/o0t1ng0o/TW-GAN (accessed on 2023-12-16), Karlsson and Hardarson, (2022) [^7] https://github.com/robert-karlsson/av-segmentation (accessed on 2023-12-16), and Galdran et al., (2022)  [^8] https://github.com/agaldran/lwnet (accessed on 2023-12-16). In order to standardize the evaluation criteria, we did not perform any post-processing on the segmentation maps produced by these methods. Following previous work Galdran et al., (2019); Chen et al., (2022); Galdran et al., (2022), models were trained and tested separately on RITE and HRF, while LES-AV was used for cross-dataset evaluation (trained on RITE, tested on LES-AV).
@@ 5.3 Evaluation metrics
Segmentation performance was evaluated using receiver operating characteristic (ROC) curves, precision-recall (PR) curves, and one-versus-all classification metrics (sensitivity, specificity, and accuracy) for each structure of interest. We calculated the area under the curve (AUC) value to summarize the information from the curves. The use of PR curves along with ROC curves was motivated by their greater sensitivity to imbalanced classes Davis and Goadrich, (2006); Saito and Rehmsmeier, (2015), as encountered here with arteries, veins, and background (see Table 1). For arteries and veins, only pixels within the ROI and excluding uncertain vessels and crossings were considered, aligning with common practices in the literature Girard et al., (2019); Hemelings et al., (2019); Galdran et al., (2022); Karlsson and Hardarson, (2022). This ensures a fair comparison with other works that typically disregard crossings during evaluation. In each case, the positive class is the structure of interest, and the negative class is everything else within the ROI.
A/V and BV/BG classification performances were evaluated using one-versus-one evaluation metrics (sensitivity, specificity, and accuracy), considering arteries and BV as the positive class, respectively. Only vessel pixels excluding crossings and uncertain vessels were involved in the calculation. While most prior works Girard et al., (2019); Hemelings et al., (2019); Galdran et al., (2022); Karlsson and Hardarson, (2022) only consider the intersection between predicted and GT vessels, this approach can yield misleading performance measures, especially for poor segmentations, and hinders standardized comparisons. Therefore, following recent works Morano et al., (2021); Chen et al., (2022), we also report classification performance for all vessel pixels in the GT, including those that were not detected as such by the models. In our state-of-the-art comparison, we report the classification performance for both scenarios, explicitly specifying the evaluation criteria used.
Additionally, we assessed the topological connectivity of A/V segmentation maps using infeasible (INF) and correct (COR) path percentages, as introduced in Araújo et al., (2019) [^9] https://github.com/rjtaraujo/dvae-refiner (accessed on 2023-12-20) and adopted by Chen et al., (2022). These metrics involve randomly sampling paths from both GT and generated masks, classifying them as infeasible if they are absent in the generated mask and correct if they differ by less than 10% from the GT. Higher COR and lower INF values indicate more topologically accurate segmentations.
@@ 5.4 Training and evaluation details
We employed 4-fold cross-validation on the RITE and HRF training sets, dividing each fold into 80% for training and 20% for validation. The Adam optimizer Kingma and Ba, (2015) with a constant learning rate of $\alpha=1\times 10^{-4}$ and decay rates of $\beta_{1}=0.9$ and $\beta_{2}=0.999$ was used for training. Early stopping was applied after 200 epochs with no decrease in validation loss. The batch size is set to 1. For the state-of-the-art comparison, models with the lowest validation error among the different folds were chosen.
We maintained the original splits of RITE and HRF for training and testing. RITE images were used at full resolution for both training and testing. HRF images were resized to 1024 pixels wide for training and testing, but predicted segmentation maps were then upsampled to the original resolution for the evaluation, following Galdran et al., (2022). Similarly, for LES-AV, we predicted at full resolution by feeding the model trained on RITE with LES-AV images resized to 576 pixels wide and then upsampling the predictions.
All images underwent offline preprocessing including global contrast enhancement and local intensity normalization, following Morano et al., (2021). Online data augmentation with color/intensity variations, affine transformations, flipping, and random cutout was applied during training.
@@ 6.1 Hyperparameter search
Refer to captionRefer to captionRefer to caption
$k=2$
$k=3$
$k=4$
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
$k=5$
$k=6$
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
@@ 6.2 Ablation study
Table 3: Ablation study in RITE. W-Net: 2 stacked U-Nets without recursive refinement ($K=1$), similar to Galdran et al., (2022). RRU-Net: recursive refinement using a single U-Net module, similar to Mosinska et al., (2018). RRWNetAll: proposed approach with recursive refinement ($K=6$) for arteries, veins, and BV. RRWNet: proposed approach with recursive refinement of arteries and veins only ($K=6$). A one-tailed Wilcoxon signed-rank test was performed to compare the results of the proposed RRWNet with the results of the best or second best model for each evaluation metric. *: $p<0.05$. Best results are highlighted in bold, and second best results are underlined.
@@
Unlike the previous categories (A/V/BV segmentation and A/V classification), methods employing refinement strategies did not consistently outperform the U-Net baseline in BV/BG classification, with slight performance reductions observed in 7 out of 12 metrics. This decrease was particularly evident for RRWNetAll, which yielded a 0.2 pp lower accuracy compared to the U-Net baseline, consistently with its BV segmentation peformance. Despite the observed decrease in certain metrics for some refinement methods, RRWNet remained the best performing method in terms of both sensitivity and accuracy for BV/BG classification, with significant improvements over other methods in the latter metric. The observed discrepancy in performance between BV/BG classification and BV segmentation, where all metrics exhibited statistically significant differences, can be partially attributed to the threshold employed for binarizing the model outputs. In BV segmentation, the evaluation metrics (AUROC and AUPR) are inherently threshold-independent, circumventing this potential issue. However, classification metrics were computed using the binary segmentation maps obtained after applying a threshold of 0.5 to the predicted probability maps, following Morano et al., (2021) and Karlsson and Hardarson, (2022). While this threshold optimizes accuracy on the training set, it may not generalize optimally to the test set, potentially explaining the observed performance difference.
@@
Refer to captionRefer to captionRefer to caption
RR-UNet
RRWNetAll
RRWNet
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to captionRefer to caption
@@ 6.3.1 A/V classification and BV segmentation
RITE Girard et al., (2019) 86.3 86.6 86.5 78.4 98.1 95.7 97.2
Galdran et al., (2019) 89 90 89 94 93 93 95
Ma et al., (2019) 93.4 95.5 94.5 79.16 98.11 95.70 98.10
Hemelings et al., (2019)* 95.13 92.78 93.81 77.61 98.74 96.08 88.17
Kang et al., (2020) 88.63 92.72 90.81 - - - -
Morano et al., (2021) 87.47 90.89 89.24 79.12 98.65 96.16 98.33
Galdran et al., (2022)* 88.86 96.04 92.76 83.05 98.19 96.29 98.47
Hatamizadeh et al., (2022) 93.10 94.31 95.13 - - - -
Karlsson and Hardarson, (2022) 95.1 96.0 95.6 82.2 97.6 95.6 98.1
Chen et al., (2022) 95.38 97.20 96.34 81.51 97.81 95.75 96.29
Hu et al., (2024) 93.37 95.37 94.42 79.08 98.15 95.69 98.07
Second expert Qureshi et al., (2013) 95.80 96.82 96.37 80.38 96.83 94.76 -
@@
Galdran et al., (2022)$\ddagger$ 86.86 93.56 90.47 76.40 97.73 95.69 96.27
RRWNet (ours)normal-$\ddagger$ 94.30 95.25 94.81 86.41 96.59 95.61 97.72
HRF Galdran et al., (2019) 85 91 91 - - - -
@@
Chen et al., (2022)* 97.06 97.29 97.19 78.14 98.29 96.59 94.66
Galdran et al., (2022)* 98.10 93.17 95.35 81.19 98.12 96.70 98.55
Karlsson and Hardarson, (2022)* 97.07 96.53 96.77 86.17 97.09 96.17 98.42
Hu et al., (2024) 93.37 95.37 94.42 69.01 99.02 96.25 98.15
Second expert Hemelings et al., (2019) 97.46 97.05 97.23 93.85 98.91 98.48 -
@@
In RITE, RRWNet achieved an A/V classification accuracy of 96.66% and a BV segmentation AUROC of 98.50%. These results surpass the second-best methods, Chen et al., (2022) and Morano et al., (2021), by 0.32 pp and 0.27 pp, respectively. Furthermore, it exceed the performance of the Second Expert Qureshi et al., (2013) in terms of A/V and BV/BG classification by 0.29 pp and 1.53 pp, respectively, demonstrating human-level performance in both tasks.
The proposed RRWNet also achieved state-of-the-art performance in the LES-AV dataset, with an A/V classification accuracy of 94.81% and a BV segmentation AUROC of 97.72%. These values represent improvements of 2.62 pp and 1.63 pp over the second best performing methods, Kang et al., (2020) and Galdran et al., (2022), respectively. Notably, RRWNet was evaluated in a cross-dataset setting (trained on RITE, tested on LES-AV), while Kang et al., (2020) was evaluated in a 2-fold cross-validation setting within LES-AV. This showcases the robustness of RRWNet and its superior performance to generalize to unseen datasets.
Finally, in HRF, RRWNet achieved once again state-of-the-art performance, with an A/V classification Accuracy of 97.83% (+0.64 pp over Chen et al., (2022)) and BV segmentation AUROC of 98.57% (+0.35 pp over Galdran et al., (2022)).

### v3

@@ 5.1 Datasets
Experiments were performed on the three publicly available datasets containing color fundus images with corresponding A/V annotations: RITE Hu et al., [2013], LES-AV Orlando et al., [2018], and HRF Budai et al., [2013]. Fig. 5 shows some examples of color fundus images and their corresponding GT segmentation maps from the three datasets, while Table 1 provides an overview of the distribution of samples (pixels) across the different classes (background, artery, vein, crossing, uncertain) within each dataset. Further details on the datasets are provided below.
@@
The RITE [^1] https://medicine.uiowa.edu/eye/rite-dataset (accessed on 2023-12-15) dataset Hu et al., [2013] (also known as AV-DRIVE or DRIVE-AV in the literature) is an extension of the Digital Retinal Images for Vessel Extraction (DRIVE) dataset Staal et al., [2004]. While DRIVE focuses on vessel segmentation, RITE incorporates additional manual GT segmentation maps for classifying arteries and veins. The dataset comprises the same 40 retinography of DRIVE (20 for training and 20 for testing). These images originate from 33 healthy patients and 7 patients with mild signs of DR. They are all centered on the macula and have a resolution of $768\times 584$ pixels, with a circular region of interest (ROI). RITE includes the original blood vessel segmentation maps from DRIVE, along with the pixel-level classification of these vessels into arteries, veins, crossings, and uncertain classes Hu et al., [2013]. Crossings indicate areas where a vein and an artery overlap. The “uncertain” class is used for those vessels whose classification the experts have not been able to determine. Alternative A/V classification annotations for the DRIVE dataset were proposed by Qureshi et al., [2013]. They manually segmented and classified the blood vessels into arteries, veins, and uncertain from the raw retinography images, independent of the existing DRIVE vessel segmentations. For artery-vein crossings, the classification was assigned based on the vessel closest to the surface of the retina. These alternative labels, henceforth referred to as Qureshi et al., were used in this work as additional “second expert” annotations.
@@
This dataset [^2] https://figshare.com/articles/dataset/LES-AV_dataset/11857698 (accessed on 2023-12-16) Orlando et al., [2018] consists of 22 retinography images, originating from both healthy patients (11) and patients with signs of glaucoma (11). Unlike RITE, LES-AV does not have a predefined split for training and testing. The images are centered on the optic disc and have a resolution of $1620\times 1444$ pixels (except one image with a resolution of $2196\times 1958$ pixels) and a circular ROI. Similarly to RITE, LES-AV includes GT segmentation maps for blood vessels, classified into arteries, veins, crossings, and uncertain regions. In this work, we employ LES-AV as an external dataset for cross-dataset evaluations to assess the generalization capabilities of the proposed method.
@@
The HRF dataset [^3] https://www5.cs.fau.de/research/data/fundus-images/ (accessed on 2023-12-16) Budai et al., [2013] is a collection of 45 high-resolution retinography images ($3504\times 2336$ pixels). The images are categorized into three groups: 15 images from healthy individuals, 15 from patients with diabetic retinopathy (DR), and 15 from glaucomatous patients. A/V classification annotations were provided by two different sources. Hemelings et al., [2019] [^4] https://github.com/rubenhx/av-segmentation (accessed on 2023-12-16) provided the original annotations used for the dataset. Then, Chen et al., [2022] [^5] https://github.com/o0t1ng0o/TW-GAN (accessed on 2023-12-16) introduced novel manual annotations to address inconsistencies in the annotations of Hemelings et al., [2019]. The annotation procedures of these two works differ slightly in handling artery-vein crossings. While Chen et al., [2022] label crossings as a separate class (consistent with RITE and LES-AV), Hemelings et al., [2019] assign them to the uppermost vessel if known, or to the uncertain class otherwise. In this work, we primarily use the Chen et al., [2022] annotations for training and testing, while we use the Hemelings et al., [2019] annotations as “second expert” annotations. Following previous works Hemelings et al., [2019], Karlsson and Hardarson, [2022], we use the first five images in each category for testing and the remainder for training.
@@ 5.2 Experiments
We performed an ablation study to evaluate the individual and combined impact of the proposed RR module and the recursive refinement (multiple RR applications). For this end, we compared the following models: (1) U-Net: A standard U-Net architecture Ronneberger et al., [2015]. This is equivalent to use only the Base subnetwork of RRWNet. (2) W-Net: Our proposed approach without recursive refinement (i.e., $K=1$). This isolates the effect of the recursive refinement. (3) RRU-Net: An architecture similar to Mosinska et al., [2018], that employs recursive refinement with a single encoder-decoder. This is equivalent to use only the RR subnetwork of RRWNet in a recursive manner while providing the input image. (4) RRWNetAll: Our full model (Base and RR subnetworks) refining all segmentation maps (arteries, veins, and BV). (5) RRWNet: Our full model refining only the A/V segmentation maps. All experiments were conducted on the RITE dataset. Statistical significance was assessed using the one-tailed Wilcoxon signed-rank test.
@@
We compared RRWNet with several state-of-the-art methods for A/V segmentation and classification on RITE, LES-AV and HRF datasets: Girard et al., [2019], Galdran et al., [2019], Ma et al., [2019], Hemelings et al., [2019], Kang et al., [2020], Morano et al., [2021], Chen et al., [2022], Galdran et al., [2022], Hatamizadeh et al., [2022], Karlsson and Hardarson, [2022], Hu et al., [2024]. To comprehensively evaluate our approach, we used different evaluation protocols. In the first place, we employed the standard evaluation protocol in the field, focusing on the A/V classification accuracy for the intersection of predicted and GT vessels. Additionally, we performed a more in depth comparison with recent state-of-the-art methods that provided source code or continuous predicted segmentation maps for any of the datasets. In particular, following Morano et al., [2021] and Chen et al., [2022], we compared the methods in terms of A/V classification performance for all vessel pixels in the GT, as well as as A/V segmentation performance using and several threshold-independent and topological metrics. The methods included in the comparison were Morano et al., [2021], Chen et al., [2022] [^6] https://github.com/o0t1ng0o/TW-GAN (accessed on 2023-12-16), Karlsson and Hardarson, [2022] [^7] https://github.com/robert-karlsson/av-segmentation (accessed on 2023-12-16), and Galdran et al., [2022]  [^8] https://github.com/agaldran/lwnet (accessed on 2023-12-16). In order to standardize the evaluation criteria, we did not perform any post-processing on the segmentation maps produced by these methods. Following previous work Galdran et al., [2019], Chen et al., [2022], Galdran et al., [2022], models were trained and tested separately on RITE and HRF, while LES-AV was used for cross-dataset evaluation (trained on RITE, tested on LES-AV).
@@ 5.3 Evaluation metrics
Segmentation performance was evaluated using receiver operating characteristic (ROC) curves, precision-recall (PR) curves, and one-versus-all classification metrics (sensitivity, specificity, and accuracy) for each structure of interest. We calculated the area under the curve (AUC) value to summarize the information from the curves. The use of PR curves along with ROC curves was motivated by their greater sensitivity to imbalanced classes Davis and Goadrich, [2006], Saito and Rehmsmeier, [2015], as encountered here with arteries, veins, and background (see Table 1). For arteries and veins, only pixels within the ROI and excluding uncertain vessels and crossings were considered, aligning with common practices in the literature Girard et al., [2019], Hemelings et al., [2019], Galdran et al., [2022], Karlsson and Hardarson, [2022]. This ensures a fair comparison with other works that typically disregard crossings during evaluation. In each case, the positive class is the structure of interest, and the negative class is everything else within the ROI.
A/V and BV/BG classification performances were evaluated using one-versus-one evaluation metrics (sensitivity, specificity, and accuracy), considering arteries and BV as the positive class, respectively. Only vessel pixels excluding crossings and uncertain vessels were involved in the calculation. While most prior works Girard et al., [2019], Hemelings et al., [2019], Galdran et al., [2022], Karlsson and Hardarson, [2022] only consider the intersection between predicted and GT vessels, this approach can yield misleading performance measures, especially for poor segmentations, and hinders standardized comparisons. Therefore, following recent works Morano et al., [2021], Chen et al., [2022], we also report classification performance for all vessel pixels in the GT, including those that were not detected as such by the models. In our state-of-the-art comparison, we report the classification performance for both scenarios, explicitly specifying the evaluation criteria used.
Additionally, we assessed the topological connectivity of A/V segmentation maps using infeasible (INF) and correct (COR) path percentages, as introduced in Araújo et al., [2019] [^9] https://github.com/rjtaraujo/dvae-refiner (accessed on 2023-12-20) and adopted by Chen et al., [2022]. These metrics involve randomly sampling paths from both GT and generated masks, classifying them as infeasible if they are absent in the generated mask and correct if they differ by less than 10% from the GT. Higher COR and lower INF values indicate more topologically accurate segmentations.
@@ 5.4 Training and evaluation details
We employed 4-fold cross-validation on the RITE and HRF training sets, dividing each fold into 80% for training and 20% for validation. The Adam optimizer Kingma and Ba, [2015] with a constant learning rate of $\alpha=1\times 10^{-4}$ and decay rates of $\beta_{1}=0.9$ and $\beta_{2}=0.999$ was used for training. Early stopping was applied after 200 epochs with no decrease in validation loss. The batch size is set to 1. For the state-of-the-art comparison, models with the lowest validation error among the different folds were chosen.
We maintained the original splits of RITE and HRF for training and testing. RITE images were used at full resolution for both training and testing. HRF images were resized to 1024 pixels wide for training and testing, but predicted segmentation maps were then upsampled to the original resolution for the evaluation, following Galdran et al., [2022]. Similarly, for LES-AV, we predicted at full resolution by feeding the model trained on RITE with LES-AV images resized to 576 pixels wide and then upsampling the predictions.
All images underwent offline preprocessing including global contrast enhancement and local intensity normalization, following Morano et al., [2021]. Online data augmentation with color/intensity variations, affine transformations, flipping, and random cutout was applied during training.
@@ 6.1 Hyperparameter search
Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption
$k=2$ $k=3$ $k=4$
@@
Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption
$k=5$ $k=6$
Refer to caption Refer to caption
@@ 6.2 Ablation study
Table 3: Ablation study in RITE. W-Net: 2 stacked U-Nets without recursive refinement ($K=1$), similar to Galdran et al., [2022]. RRU-Net: recursive refinement using a single U-Net module, similar to Mosinska et al., [2018]. RRWNetAll: proposed approach with recursive refinement ($K=6$) for arteries, veins, and BV. RRWNet: proposed approach with recursive refinement of arteries and veins only ($K=6$). A one-tailed Wilcoxon signed-rank test was performed to compare the results of the proposed RRWNet with the results of the best or second best model for each evaluation metric. *: $p<0.05$. Best results are highlighted in bold, and second best results are underlined.
@@
Unlike the previous categories (A/V/BV segmentation and A/V classification), methods employing refinement strategies did not consistently outperform the U-Net baseline in BV/BG classification, with slight performance reductions observed in 7 out of 12 metrics. This decrease was particularly evident for RRWNetAll, which yielded a 0.2 pp lower accuracy compared to the U-Net baseline, consistently with its BV segmentation peformance. Despite the observed decrease in certain metrics for some refinement methods, RRWNet remained the best performing method in terms of both sensitivity and accuracy for BV/BG classification, with significant improvements over other methods in the latter metric. The observed discrepancy in performance between BV/BG classification and BV segmentation, where all metrics exhibited statistically significant differences, can be partially attributed to the threshold employed for binarizing the model outputs. In BV segmentation, the evaluation metrics (AUROC and AUPR) are inherently threshold-independent, circumventing this potential issue. However, classification metrics were computed using the binary segmentation maps obtained after applying a threshold of 0.5 to the predicted probability maps, following Morano et al., [2021] and Karlsson and Hardarson, [2022]. While this threshold optimizes accuracy on the training set, it may not generalize optimally to the test set, potentially explaining the observed performance difference.
@@
Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption
RR-UNet RRWNetAll RRWNet
@@
Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption Refer to captionRefer to captionRefer to caption
@@ 6.3.1 A/V classification and BV segmentation
RITE Girard et al., [2019] 86.3 86.6 86.5 78.4 98.1 95.7 97.2
Galdran et al., [2019] 89 90 89 94 93 93 95
Ma et al., [2019] 93.4 95.5 94.5 79.16 98.11 95.70 98.10
Hemelings et al., [2019]* 95.13 92.78 93.81 77.61 98.74 96.08 88.17
Kang et al., [2020] 88.63 92.72 90.81 - - - -
Morano et al., [2021] 87.47 90.89 89.24 79.12 98.65 96.16 98.33
Galdran et al., [2022]* 88.86 96.04 92.76 83.05 98.19 96.29 98.47
Hatamizadeh et al., [2022] 93.10 94.31 95.13 - - - -
Karlsson and Hardarson, [2022] 95.1 96.0 95.6 82.2 97.6 95.6 98.1
Chen et al., [2022] 95.38 97.20 96.34 81.51 97.81 95.75 96.29
Hu et al., [2024] 93.37 95.37 94.42 79.08 98.15 95.69 98.07
Second expert Qureshi et al., [2013] 95.80 96.82 96.37 80.38 96.83 94.76 -
@@
Galdran et al., [2022]$\ddagger$ 86.86 93.56 90.47 76.40 97.73 95.69 96.27
RRWNet (ours)$\ddagger$ 94.30 95.25 94.81 86.41 96.59 95.61 97.72
HRF Galdran et al., [2019] 85 91 91 - - - -
@@
Chen et al., [2022]* 97.06 97.29 97.19 78.14 98.29 96.59 94.66
Galdran et al., [2022]* 98.10 93.17 95.35 81.19 98.12 96.70 98.55
Karlsson and Hardarson, [2022]* 97.07 96.53 96.77 86.17 97.09 96.17 98.42
Hu et al., [2024] 93.37 95.37 94.42 69.01 99.02 96.25 98.15
Second expert Hemelings et al., [2019] 97.46 97.05 97.23 93.85 98.91 98.48 -
@@
In RITE, RRWNet achieved an A/V classification accuracy of 96.66% and a BV segmentation AUROC of 98.50%. These results surpass the second-best methods, Chen et al., [2022] and Morano et al., [2021], by 0.32 pp and 0.27 pp, respectively. Furthermore, it exceed the performance of the Second Expert Qureshi et al., [2013] in terms of A/V and BV/BG classification by 0.29 pp and 1.53 pp, respectively, demonstrating human-level performance in both tasks.
The proposed RRWNet also achieved state-of-the-art performance in the LES-AV dataset, with an A/V classification accuracy of 94.81% and a BV segmentation AUROC of 97.72%. These values represent improvements of 2.62 pp and 1.63 pp over the second best performing methods, Kang et al., [2020] and Galdran et al., [2022], respectively. Notably, RRWNet was evaluated in a cross-dataset setting (trained on RITE, tested on LES-AV), while Kang et al., [2020] was evaluated in a 2-fold cross-validation setting within LES-AV. This showcases the robustness of RRWNet and its superior performance to generalize to unseen datasets.
Finally, in HRF, RRWNet achieved once again state-of-the-art performance, with an A/V classification Accuracy of 97.83% (+0.64 pp over Chen et al., [2022]) and BV segmentation AUROC of 98.57% (+0.35 pp over Galdran et al., [2022]).